OpenStudy (anonymous):

Need help A model airplane is shot up from a platform 1 foot above the ground with an initial upward velocity of 56 feet per second. The height of the airplane above ground after t seconds is given by the equation h=-16t^2+56t+1, where h is the height of the airplane in feet and t is the time in seconds after it is launched. Approximately how long does it take the airplane to reach its maximum height? A. 0.3 seconds B. 1.8 seconds C. 3.5 seconds D. 6.9 seconds

OpenStudy (anonymous):

I say it's B but no sure

OpenStudy (nurali):

v = Vi - 32 t at top v = 0 0 = 56 - 32 t t = 1.75 ≈ 1.8 seconds